3m/4+m/2=-5 solve for m
Found 2 solutions by AnlytcPhil, JimMarshall:
Answer by AnlytcPhil(1806)


>>3m/4+m/2=-5 solve for m<<
`
` ` ` ` `3m/4 + m/2 = -5
`
` Multiply through by LCD = 4
`
` ` `4·3m/4 + 4·m/2 = 4·-5
`
` Cancel the 4's in the first term, get 3m
` Divide the 4 by 2 in the 2nd term, get 2m
` Multiply the 4 by -5, get -20
`
` ` ` ` ` ` 3m + 2m = -20
`
` Combine 3m and 2m, get 5m
`
` ` ` ` ` ` ` ` `5m = -20
`
` Divide both sides by 5
`
` ` ` ` ` ` ` `5m/5 = -20/5
`
` ` ` ` ` ` ` ` ` m = -4
`
Edwin

Answer by JimMarshall(82)

3m / 4 + m/2 = -5
==> (3m +2m )/4 = -5 ( LCM of 2 & 4 = 4 )
==> 5m=-20
==> m=-4
ANS